OpenStudy (anonymous):

Express 0.00005 s in µs.

OpenStudy (theeric):

This isn't so bad when you get the concept! Metric is nice! When you get 1,000 of one unit, it's 1 of another! Or, if you have \(\dfrac 1{1,000}\) of one unit, you have one of another. So, here's the order: 1,000 nanoseconds (ns) = 1 microsecond (\(\mu\)s) 1,000 microseconds (\(\mu\)s) = 1 millisecond (ms) 1,000 milliseconds (ms) = 1 second (s) But you don't have to take it in steps. For example, \(\rm1s = 1,000ms = 1,000,000 \mu s = 1,000,000,000ns\) This comes from "unit conversion." Please ask about that if you need! It's easier to say that \(\rm 1s = 10^3ms=10^6\mu s=10^9ns\) It's a lot to take in, but you can see that it's still increasing by a factor of 1,000 each time... Since \(\rm1s=10^6\mu s\), you convert like before (I didn't show it, but ask if you want to see it) \(\rm0.00005s*1=0.00005s*\dfrac{10^6\mu s}{1s}=?\) The answer will be in \(\mu\rm s\), see? So, that number, 0.00005 will be 1,000,000 times larger. So, move the decimal 6 places to the right. Each place you move to the right is just like multiplying by 10! So multiply by 10 a total of 6 times. Move the decimal place to the right 6 digits.